OpenStudy (anonymous):

3x/7 +7/5 = 12 please show work so I may understand how to do this.. thank you so much

OpenStudy (anonymous):

how did you get the 53?

OpenStudy (diyadiya):

12-7/5= (60-7)/5=53/5

OpenStudy (anonymous):

thank you..